OpenStudy (anonymous):

What is the area of a rectanlge with the vertices (4,6), (0,3), (3,-1) and (7,2)? a)10 b) 15 c) 20 d) 25

OpenStudy (anonymous):

Apply determinant

OpenStudy (anonymous):

I dont understand determinant

OpenStudy (aripotta):

use the distance formula

OpenStudy (aripotta):

the distance formula will give you the lengths of the sides. luckily, you only have to do that twice, not 4 times, because rectangles have congruent opposite sides

OpenStudy (anonymous):

i got 10, both we're 5 & 5

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }|\left[\begin{matrix}4 & 6 \\ 0 & 3\end{matrix}\right]+\left[\begin{matrix}0 & 3 \\ 3 & -1\end{matrix}\right]+\left[\begin{matrix}3 & -1 \\ 7 & 2\end{matrix}\right]\left[\begin{matrix}7 & 2 \\ 4 & 6\end{matrix}\right]|=25\]

OpenStudy (aripotta):

so 5 x 5 = 25

OpenStudy (aripotta):

and that's our area